Additional Information: During the year, Sunland sold equipment with an original cost of $143,820 and accumulated depreciation of $111,860 and purchased new equipment for $255,680. Prepare a statement of cash flows for the year ending December 31, 2020.Sunland Enterprises Balance Sheet and Income Statement Data December 31, 2020 December 31, 2019 Current Assets: Cash $143,820
Answer #1

Answer:

$154,536 Sunland Enterprises Statement of Cash Flows For the year ended December 31, 2020 Cash flow from Operating Activities

Workings;

Dividend = Retained earnings beg.balance + Net Income - Retained earnings ending balance

= 319600 + 154536 - 351560

= $122,576

Sale value of equipment = Cost - Accumulated depreciation - Loss on sale =143820 - 111860 - 11280 = $20,680